# An X-Ray High-Frequency QPO in NGC 1365

### v1

@@ 2.2 Time-averaged Spectra Analysis
To study the spectral properties of NGC 1365 more accurately, XSPEC (Version 12.13.1n, Arnaud (1996)) was employed for the spectral analysis. The energy range we used is data from 1.0 – 10.0 keV. As a starting point, An initial simple model, consisting of three distinct components related to AGN, was utilized: $TBabs\times(powerlaw+pexrav+apec)$ (Nardini et al., 2015), to simultaneously fit the spectra from three EPIC cameras. In this model, TBabs is the Tuebingen-Boulder ISM absorption model, representing the galactic absorption of NGC 1365, fixed at $N_{\rm H}=1.34\times 10^{20}$ cm-2 (according to Kalberla et al. (2005)). For the part reflected by the irradiated material, we used a pexrav model (Magdziarz & Zdziarski, 1995) to describe it, which has solar abundance, an inclination angle of 45, and variable intensity. It also includes a soft X-ray power law, interpreted as a dim, scattered AGN continuum. At the same time, to check whether a part of the emission line can be attributed to optically thin and collisionally ionized plasma, which is an important component of the XMM–Newton/RGS spectrum (Guainazzi et al., 2009), we added an apec thermal component with solar abundance (Smith et al., 2001). We used the above model as the initial model for fitting.
@@
The fitting result of this model shows a significant negative deviation at 6.7 – 8.3 keV, a phenomenon that has been discussed in several previous papers (Risaliti et al., 2005, 2009). This is because there are four absorption lines in the X-ray spectrum of NGC 1365 in this range. Through spectral analysis, we identified these features as the $K\alpha$ and $K\beta$ lines of Fe xxv and Fe xxvi. Specifically, the corresponding energies of Fe xxv $K\alpha$ and Fe xxvi $K\beta$ are 6.697 keV and 6.966 keV, respectively, while those of Fe xxv $K\alpha$ and Fe xxvi $K\beta$ are 7.880 keV and 8.268 keV (Risaliti et al., 2005). Therefore, we multiplied two gabs models (Maitra et al., 2018; Klochkov et al., 2011) to describe these absorption lines.
@@
First, we fixed the LineE of the two gabs models at 6.832 keV and 8.074 keV, respectively, and released them after the other parameters were well fitted. In the end, the LineE of the two gabs models were fitted at 6.837 keV and 7.988 keV, respectively. The new fit (Model B) brought significant improvement, with the goodness of fit is $\chi^{2}/\nu=1.27$. However, when we checked the fitting data, we found that the fitting result of the pexrav model had a large deviation, with a negative photon index power and a reflection scaling parameter very close to zero. After discussion, it was determined that the reflection model was not suitable and was therefore removed. NGC 1365 is considered to show partially obscured X-ray absorption (Risaliti et al., 2009; Maiolino et al., 2010), so we introduced the partially obscured model TBpcf (Swain et al., 2023) for description and multiplied it with the power law component.
@@
Therefore, we obtained the final fitting model(model D): $Tbabs\times gabs\times gabs\times gabs\times gabs\times(gaussian+apec+tbpcf\times powerlaw)$. Based on the original fitting model C, we first fixed the LineE of the gaussian at 6.40 keV for fitting, then released it, and finally fitted at 6.35 keV. This model provides a good fit, with an overall statistic of $\chi^{2}/\nu=1.10$. We compared the results we obtained with the fitting data from the studies of Risaliti et al. (2005), Nardini et al. (2015) and Swain et al. (2023), and the results are somewhat consistent. The best-fitting parameters and related errors are shown in Table 2. The spectrum of the best-fitting model is displayed in the middle panel of Figure 4, the top panel shows the ld plot, and the bottom panel shows the residuals.
@@
Figure 4: XMM-Newton spectrum of NGC 1365 (ObsID 0205590301). Top and middle panel are the logarithmic data and eeuf diagram. Black, red and blue represent the data of PN, MOS1 and MOS2, respectively. The best-fitting model of model D and the residuals ([data–model]/error) are shown in lower panel.
Figure 5: Correlation between the black hole masses and their QPO frequencies. The events reported in previous works are maked with green points and the new QPO signal detected in NGC 1365 is marked with a red square. The three lines represent the relations suggested in Remillard & McClintock (2006) and Kluzniak & Abramowicz (2002). See Zhou et al. (2010, 2015) and Zhang et al. (2017, 2018) for additional information.
@@ 3 SUMMARY AND DISCUSSION
In this paper, the observational data from the XMM-Newton X-ray telescope for the Seyfert galaxy NGC 1365, obtained in January 2004, were analyzed in this research. By employing the WWZ method to construct the time-averaged power spectrum and analyzing it with the LSP method, we identified a QPO signal with a period of 4566 s and a confidence level of 3.6 $\sigma$ throughout the observation period. However, we did not detect this signal in other XMM-Newton observational data. Therefore, we believe that this signal, like the QPO signals found in Mrk 766 (Zhang et al., 2017), 1H 0707--495 (Pan et al., 2016; Zhang et al., 2018), and Swift J1644$+$57 (Zhou et al., 2015), is a transient periodic signal.
@@
Currently, the origin of QPO remains a topic of debate in the academic community. However, existing research generally believes that QPO signals may be generated by several mechanisms:Pulsating accretion near the Eddington limit;Instability of the inner accretion disk; X-ray hotspots orbiting the black hole or disk precession (according to the Bardeen-Petterson effect) (see Remillard & McClintock (2006); Li et al. (2003); Syunyaev (1973); Guilbert et al. (1983); Bardeen & Petterson (1975); Mukhopadhyay et al. (2003); Gangopadhyay et al. (2012). The related explanatory models include the resonance model (Abramowicz & Kluźniak, 2001), the relativistic precession model (Stella et al., 1999), the acoustic oscillation model (Rezzolla et al., 2003), the accretion-ejection instability in the accretion disk (Tagger & Pellat, 1999), and disk oscillations under different modes (g, c, p mode) (Wagoner, 1999).
@@
At the same time, these models also predict the relationship between $f_{\rm QPO}$ and $M_{\rm BH}$. In the $f_{\rm QPO}-M_{\rm BH}$ correlation diagram, we first display some of the previous research (Remillard & McClintock, 2006; Kluzniak & Abramowicz, 2002; Zhou et al., 2010, 2015; Pan et al., 2016), As shown in Figure 5. The QPO frequencies found in RE J1034+396 (Gierlinski et al., 2008) and 1H 0707--495 (Pan et al., 2016; Zhang et al., 2018) have been considered to be HFQPOs (Zhou et al., 2010, 2015). The QPO frequency of NGC 1365 and the mass of the central black hole are plotted on the graph, demonstrating a high level of consistency. This suggests that the discovered QPO may also be classified as HFQPO. The mass data comes from the calculation obtained by Fazeli et al. (2019): the mass of the central black hole is in the range of (5 – 10) ×$\times 10^{6}\ M_{\odot}$. The origin of QPO in black hole X-ray binaries and Seyfert galaxies is still unclear, and our results can provide more information for this research.
@@
We attempted to fit the energy spectrum ranging from 0.2 – 10.0 keV, but regardless of the model employed, we were unable to achieve a satisfactory fit. Consequently, we calculated the energy spectra for all observations of NGC 1365, and the results indicated that they exhibit minimal variance in 0.2 – 1.0 keV range. None of them exhibited QPO signals, thus leading us to believe that the processes generating QPO signals predominantly emit hard X-ray photons with energies exceeding 1.0 keV. After excluding photons without QPO signals, the signal’s confidence level was enhanced, reaching 3.9 $\sigma$. We did not observe a situation where QPOs only appeared in the high energy band in Mrk 766(Zhang et al., 2017) and 1H 0707-495(Zhang et al., 2018). This scenario likely indicates that the region where QPOs occur primarily emits high-energy X-ray radiation.
@@
The X-ray data in the 1.0 – 10.0 keV band were fitted with the apec model, primarily for the soft band, and the TBpcf model, which accounts for partial obscuration, was used in conjunction with a powerlaw to describe the hard band. The fitting results of the spectrum show that the plasma temperature reaches 0.791 keV, which is close to the results obtained in Swain et al. (2023). Also, NGC 1365 has multiple absorption lines and emission lines in the high-energy band (¿ 6.0 keV), which has been studied many times before (Risaliti et al., 2005, 2009; Nardini et al., 2015; Guainazzi et al., 2009), and these emission lines also show puzzling asymmetry (Lena et al., 2016). At the same time, we applied the same model to perform spectral fitting again for the peak and valley regions in the light curve, and compared the model fitting results of the peaks and valleys. The analysis shows that there is no significant difference between the best fitting data of the peaks and valleys, indicating that the QPO phenomenon is not caused by the spectrum. Considering the causes of other HFQPOs, we believe that the QPO discovered in this study may also be caused by the resonance mechanism.(Abramowicz & Kluźniak, 2001)
@@
At the same time, the gas around the black hole causes the absorption or reflection body to show strong heterogeneity, so the reflection intensity of the gas with the same column density measured by absorption is systematically stronger than expected, and there is partial obscuration (Risaliti et al., 2009; Maiolino et al., 2010), therefore, the $N_{\rm H}$ varies greatly under different observation conditions. Unlike standard Seyfert galaxies such as NGC 2617, the unstable behavior of NGC 1365 below 10.0 keV makes it difficult to obtain strictly consistent information from different observation times, which also leads to the lack of a unified conclusion on the classification of NGC 1365. The above complex properties make the spectrum of NGC 1365 difficult to describe simply. Therefore, more potential observations with high spatial resolution and high spectral resolution on NGC 1365 will help to study the sub-plasma scale environmental properties of active galaxies in depth.
Acknowledgements
@@
This work is supported by the National Natural Science Foundation of China under grants 12203029, 12373030, 12233002, U2031205, 2021YFA0718500 and 2023AFB577. The data used in this paper were obtained from the XMM-Newton, and we express our gratitude.

### v2

@@ 2.2 Time-averaged spectral analysis
To study the spectral properties of NGC 1365 more accurately, XSPEC (Version 12.13.1n, Arnaud 1996) was employed for the spectral analysis. The energy range we used is from 1.0 to 10.0 keV. As a starting point, we used a simple initial model consisting of three distinct components related to AGNs: $TBabs\times(powerlaw+pexrav+apec)$ (Nardini et al., 2015), to simultaneously fit the spectra from three EPIC cameras. In this model, TBabs is the Tuebingen-Boulder interstellar medium absorption model, representing the galactic absorption of NGC 1365, fixed at $N_{\rm H}=1.34\times 10^{20}$ cm-2 (Kalberla et al., 2005). We used a pexrav model (Magdziarz & Zdziarski, 1995) to describe the part reflected by the irradiated material; this model has solar abundance, an inclination angle of 45, and variable intensity. It also includes a soft X-ray power law, interpreted as a dim, scattered AGN continuum. At the same time, to check whether or not a part of the emission line can be attributed to optically thin and collisionally ionized plasma, which is an important component of the XMM-Newton/RGS spectrum (Guainazzi et al., 2009), we added an apec thermal component with solar abundance (Smith et al., 2001). We used the above model as the initial model for fitting.
@@
The fitting result of this model shows a significant negative deviation at 6.7 – 8.3 keV, a phenomenon that has been discussed in several previous papers (Risaliti et al., 2005, 2009). This is because there are four absorption lines in the X-ray spectrum of NGC 1365 in this range. Through spectral analysis, we identified these features as the $K\alpha$ and $K\beta$ lines of Fe xxv and Fe xxvi. Specifically, the corresponding energies of Fe xxv $K\alpha$ and Fe xxvi $K\beta$ are 6.697 keV and 6.966 keV, respectively, while those of Fe xxv $K\alpha$ and Fe xxvi $K\beta$ are 7.880 keV and 8.268 keV (Risaliti et al., 2005). Therefore, we multiplied two gabs models (Maitra et al., 2018; Klochkov et al., 2011) to describe these absorption features.
@@
First, we fixed the LineE of the two gabs models at 6.832 keV and 8.074 keV, respectively, and released them after the other parameters were well fitted. In the end, the LineE of the two gabs models were fitted at 6.837 keV and 7.988 keV, respectively. The new fit (Model B) brought significant improvement, with a goodness of fit of $\chi^{2}/\nu=1.27$. However, when we checked the fitting data, we found that the fitting result of the pexrav model had a large deviation, with a negative photon index power and a reflection scaling parameter very close to zero. After discussion, it was determined that the reflection model was not suitable and was therefore removed. NGC 1365 is considered to show partially obscured X-ray absorption (Risaliti et al., 2009; Maiolino et al., 2010), and so we introduced the partially obscured model TBpcf (Swain et al., 2023) to describe the absorption and multiplied it with the power-law component.
@@
Therefore, we obtained the final fitting model (Model D): $Tbabs\times gabs\times gabs\times gabs\times gabs\times(gaussian+apec+tbpcf\times powerlaw)$. Based on the original fitting model C, we first fixed the LineE of the Gaussian at 6.40 keV for fitting, then released it, and finally fitted at 6.35 keV. This model provides a good fit, with an overall statistic of $\chi^{2}/\nu=1.10$. We compared the results we obtained with the fitting data from the studies of Risaliti et al. (2005), Nardini et al. (2015), and Swain et al. (2023), and the results are somewhat consistent. The best-fitting parameters and related errors are shown in Table 2. The spectrum of the best-fitting model is displayed in the middle panel of Figure 4; the top panel shows the logarithmic data plot and the bottom panel shows the residuals.
@@
Figure 4: XMM-Newton spectrum of NGC 1365 (ObsID 0205590301). Top and middle panel are the logarithmic data and eeuf diagram. Black, red, and blue represent the data of PN, MOS1, and MOS2, respectively. The best-fitting model of Model D and the residuals ([data–model]/error) are shown in the lower panel.
Figure 5: Correlation between source mass and QPO frequency. The yellow points in the upper left corner represent HF QPO events observed in neutron stars, while the purple points correspond to HF QPO events in Galactic microquasar BHs. The remaining points denote HF QPO events reported in supermassive BHs. The newly detected QPO signal in NGC 1365 is marked with a red square. The three gray lines represent the relationships suggested by Remillard & McClintock (2006) and Kluzniak & Abramowicz (2002). The blue line is based on the maximum allowed orbital frequency proposed by Šrámková et al. (2021), indicating that QPO signals with frequencies above the innermost stable circular orbit would not exist in the region above and to the right of this line. For further information, please refer to the studies by Zhou et al. (2010, 2015) and Zhang et al. (2017, 2018).
@@ 3 Summary and discussion
In this study, we analyzed observational data from the XMM-Newton X-ray telescope for the Seyfert galaxy NGC 1365 obtained in January 2004. By employing the WWZ method to construct the time-averaged power spectrum and analyzing it with the LSP method, we identified a QPO signal with a period of 4566 s and a confidence level of 3.6 $\sigma$ throughout the observation period. However, we did not detect this signal in other XMM-Newton observational data. Therefore, we believe that this signal, like the QPO signals found in Mrk 766 (Zhang et al., 2017), 1H 0707--495 (Pan et al., 2016; Zhang et al., 2018), and Swift J1644$+$57 (Zhou et al., 2015), is a transient periodic signal.
@@
At present, the origin of QPOs remains a topic of debate. However, existing research generally suggests that QPO signals may be generated by several mechanisms, such as pulsating accretion near the Eddington limit, instabilities of the inner accretion disk, X-ray hotspots orbiting the BH, and disk oscillations and precession (see Syunyaev 1973; Bardeen & Petterson 1975; Guilbert et al. 1983; Mukhopadhyay et al. 2003; Li et al. 2003; Remillard & McClintock 2006; Gangopadhyay et al. 2012). The detailed explanatory models include the resonance model (Abramowicz & Kluźniak, 2001), the relativistic precession model (Stella et al., 1999), the acoustic oscillation model (Rezzolla et al., 2003), the accretion–ejection instability model (Tagger & Pellat, 1999), and disco-seismic models assuming thin disk oscillations (Wagoner, 1999).
@@
At the same time, these models also predict the relationship between $f_{\rm QPO}$ and $M_{\rm BH}$. The original diagram and associated lines were first proposed, drawn, and discussed in the works of Abramowicz et al. (2004) and Török (2005). In the $f_{\rm QPO}-M_{\rm BH}$ correlation diagram, we first display some of the previous research (Kluzniak & Abramowicz, 2002; Remillard & McClintock, 2006; Zhou et al., 2010, 2015; Pan et al., 2016), as shown in Figure 5. The QPO frequencies found in RE J1034+396 (Gierliński et al., 2008) and 1H 0707--495 (Pan et al., 2016; Zhang et al., 2018) have been considered to be HF QPOs. The QPO frequency of NGC 1365 and the mass of the central BH are plotted on the graph, demonstrating a high level of consistency. This suggests that the discovered QPO may also be classified as HF QPO. The mass data come from the calculation obtained by Fazeli et al. (2019): the mass of the central BH is in the range of (5 – 10) ×$\times 10^{6}\ M_{\odot}$. The origin of QPO in BH X-ray binaries and Seyfert galaxies is still unclear, and our results can provide more information for this research.
@@
We attempted to fit the energy spectrum ranging from 0.2 to 10.0 keV, but regardless of the model employed, we were unable to achieve a satisfactory fit. Consequently, we calculated the energy spectra for all observations of NGC 1365, and the results indicate that they exhibit minimal variance in the 0.2 – 1.0 keV range. None of them exhibit QPO signals, which leads us to believe that the processes generating QPO signals predominantly emit hard X-ray photons with energies exceeding 1.0 keV. After excluding photons without QPO signals, the signal’s confidence level was enhanced, reaching 3.9 $\sigma$. We did not observe a situation where QPOs only appeared in the high-energy band in Mrk 766 (Zhang et al., 2017) and 1H 0707-495 (Zhang et al., 2018). This scenario likely indicates that the region where QPOs occur primarily emits high-energy X-ray radiation.
@@
The X-ray data in the 1.0 – 10.0 keV band were fitted with the apec model primarily for the soft band, and the TBpcf model — which accounts for partial obscuration — was used in conjunction with a power law to describe the hard band. The fitting results of the spectrum show that the plasma temperature reaches 0.791 keV, which is close to the results obtained in Swain et al. (2023). Also, NGC 1365 has multiple absorption lines and emission lines in the high-energy band (¿ 6.0 keV), which has been studied many times before (Risaliti et al., 2005, 2009; Nardini et al., 2015; Guainazzi et al., 2009), and these emission lines also show puzzling asymmetry (Lena et al., 2016). At the same time, we applied the same model to perform spectral fitting again for the peak and valley regions in the light curve, and compared the model fitting results of the peaks and valleys. The analysis shows that there is no significant difference between the best-fitting data of the peaks and valleys, indicating that the QPO phenomenon is not caused by the spectrum. Considering the causes of other HF QPOs (Abramowicz & Kluźniak, 2001), we believe that the QPO discovered in this study may also be caused by the resonance mechanism.
@@
At the same time, the gas around the BH causes the absorption or reflection body to show strong heterogeneity, and so the reflection intensity of the gas with the same column density measured by absorption is systematically stronger than expected, and there is partial obscuration (Risaliti et al., 2009; Maiolino et al., 2010); therefore, the $N_{\rm H}$ varies greatly under different observation conditions. Unlike standard Seyfert galaxies, such as NGC 2617, the unstable behavior of NGC 1365 below 10.0 keV makes it difficult to obtain strictly consistent information from different observation times, which also means we cannot come to a unified conclusion on the classification of NGC 1365. The above complex properties mean that the spectrum of NGC 1365 cannot be summed up in a simple description. Therefore, more observations of NGC 1365 with high spatial resolution and high spectral resolution would help to study the subplasma-scale environmental properties of active galaxies in depth.
Acknowledgements.
@@
This work is supported by the National Natural Science Foundation of China under grants 12403052, 12203029, 12373030, 12233002, U2031205, 2021YFA0718500 and 2023AFB577. The data used in this paper were obtained from the XMM-Newton, and we express our gratitude.